OpenStudy (kg1975):

Plz Help 2 1/2+(3/8−1/8)

OpenStudy (kg1975):

@andrejrusa12

OpenStudy (kingnoob):

Do you mean this? \(\large \frac{1}{2} - 3(\frac{1}{2} +1) \)?

OpenStudy (kg1975):

2 power at the end and yes

OpenStudy (kingnoob):

sooo... \(\large \frac{1}{2}-3(\frac{1}{2} +1)^2 \)?

OpenStudy (kg1975):

yea idk how to write it like that but yes thats it

OpenStudy (kingnoob):

Follow "PEMDAS" So first do what's in parentheses, which is \(\large (\frac{1}{2}+1) \). Do you know how to add them?

OpenStudy (kg1975):

nope. idk how to add it like that

OpenStudy (kg1975):

im guessing 1 1/2

OpenStudy (kingnoob):

Yess, but it would be easier to handle as a fraction.. to add them just convert the 1 into a fraction. 1 is the same as \(\large \frac{2}{2} \) So \(\large \frac{1}{2}+\frac{2}{2} \)

OpenStudy (kg1975):

3/2?

OpenStudy (kingnoob):

Yess, now, do you know how to do exponents? Now we have to do the exponents which is \(\large \frac{1}{2}-3\color{red}{(\frac{3}{2})^2} \)

OpenStudy (kingnoob):

\(\large (\frac{3}{2})^2 \) is the same as \(\large \frac{3}{2}*\frac{3}{2} \) so just multiply them

OpenStudy (kg1975):

i got no clue

OpenStudy (kg1975):

1/2+(3/8−1/8)2

OpenStudy (kg1975):

i figured the other 1 out can u help with this 1

OpenStudy (kingnoob):

\(\large \frac{1}{2}+(\frac{3}{8}-\frac{1}{8})^2 \) Do the parentheses first, what do you get?

OpenStudy (kg1975):

2/8

OpenStudy (kg1975):

isnt that now.... 1/2+2/8 2

OpenStudy (kg1975):

hello?

OpenStudy (kingnoob):

Right! Now if you simplify that you get 1/4, which is easier to work with: \(\large \frac{1}{2}+(\frac{1}{4})^2 \) Now doing the exponent part, \(\large \color{red}{\frac{1}{4}^2} \) \(\large \frac{1*1}{4*4} =\frac{1}{16} \) \(\color{blue}{x^2=x*x} \) Now just add them together - \(\large \frac{1}{2}+\frac{1}{16} \) Since they don't have common denominators, you'll need to find one, which would be \(\color{red}{16} \) In order to get 16 from 2, you need to multiply by a common factor of 8 - \(\large \frac{1}{2} *\frac{8}{8} =\frac{8}{16} \) Now you can add them - \(\large \frac{8}{16} +\frac{1}{16} \)

OpenStudy (kingnoob):

make sense?